Dolores ate 3 more than Dimpy. If they ate a total of 30 pies in all, how many did each eat?
Answer by stanbon(75887)

Dolores ate 3 more than Dimpy. If they ate a total of 30 pies in all, how many did each eat?
----
L = P+3
L + P = 30
----
Substitute for "L"::
P+3+P = 30
2P = 27
P = 13.5 (Dimpy consumed 13.5 pies)
----
L = P+3 = 16.5 (Dolores consumed 16.5 pies)
-------------------
Cheers,
Stan H.
---------------------